OpenStudy (anonymous):

simplify (3a+3b)^2-2a-2b-1

OpenStudy (anonymous):

use (a+b)^2=a2+b2+2ab simply open this square and get result.

OpenStudy (triciaal):

you can actually factor out a 3 then apply the formula shown above before grouping like terms to simplify

OpenStudy (triciaal):

(3a+3b)^2 =[3(a + b)}^2 = 3^2*(a + b)^2 =9(a^2 + 2ab + b^2) = 9a^2 + 18ab + 9b^2 so (3a+3b)^2-2a-2b-1 = 9a^2 + 18ab + 9b^2 -2a -2b -1 another approach would be to use the difference of 2 squares